OpenStudy (anonymous):

Does this sequence converge or diverge? Take the limit--- (n!)/(n + 1) n= 0 to infinity

OpenStudy (amistre64):

n! always ... ALWAYS, controls the stage when n is very large, this is approximately equivalent to n!

OpenStudy (amistre64):

do a ratio test

OpenStudy (amistre64):

or simplicity\[\frac{\frac{n*(n-1)!}{n}}{\frac{n}{n}+\frac{1}{n}}\] \[\frac{(n-1)!}{1+\frac{1}{n}}\] \[\frac{(\infty-1)!}{1+\frac{1}{\infty}}=\infty\]